OpenStudy (anonymous):

What is the eccentricity of the hyperbola given by the equation (y^2/100) - (x^2/36)= 1?

OpenStudy (anonymous):

oh no! \[\sqrt{1+\frac{b^2}{a^2}}\]

OpenStudy (anonymous):

you good from there? replace \(b=36,a=100\) get \[\sqrt{1+\frac{36}{100}}\] add the numbers inside the radical and put in simplest radical form

OpenStudy (anonymous):

THANK YOU!!